Fermi National Laboratory


Accelerator Update for January 6, 2006 - January 9, 2006

Friday January 6

The day shift began with Operations monitoring store #4581, the Recycler Ring (RR) stash at 74.1 E10, Antiproton Source (Pbar) stack at 59.8 E10, and with MiniBooNE, NuMI, and MCenter (MIPP) taking beam.

Operations put the Recycler into shot setup at 6:58 AM to stash antiprotons.

Pbar returns to stacking at 7:47 AM, beam also returns to MCenter.

The Booster Chopper (B:CHOP) tripped off at 12:59 PM, taking all Proton beam away.

Experts are working to repair it.

The Chopper was fixed by 2:17 PM, beam returns for stacking, NuMI, MiniBoone, and MIPP.

Operations put the Recycler into shot setup at 4:16 PM to stash antiprotons.

Pbar returns to stacking at 5:36 PM, beam also returns to MCenter.

Booster RF station #14 (BRF14) tripped off at 7:28 PM. Any attempt to bring up the station caused lots of sparking. Experts were called to investigate.

At 8:30 PM, BRF19 also started to cause problems.

Operations put the Recycler into shot setup at 10:07 PM to stash antiprotons.

Pbar returns to stacking at 11:09 PM, beam also returns to MCenter.

Booster intensity was low for the last half of the evening shift, due to the problems with BRF14 and BRF19. BRF14 remained off and BRF19 continued to run at reduced efficiency. Lots of tuning was required to bring the Booster intensity back up.

Saturday January 7

The midnight shift began with Operations monitoring store #4581, the Recycler Ring (RR) stash at 237.5 E10, Pbar stack at 24.3 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the Recycler into shot setup at 3:39 AM to stash antiprotons.

Pbar returns to stacking at 4:29 AM, beam also returns to MCenter.

At 5:06 AM BRF4 tripped and had to be bypassed.

Operations put the Recycler into shot setup at 9:57 AM to stash antiprotons.

Store #4581 was terminated at 10:00 AM.

At 10:19 AM, CDF and D0 both made accesses to their collision halls.

Pbar returns to stacking at 10:22 AM, beam also returns to MCenter.

CDF and D0 accesses were completed at 11:53 AM.

Tevatron (Tev) studies began at 1:27 PM.

Both of MIPP’s analyzing magnets tripped off at 1:58 PM. Operations is investigating.

The analyzing magnets were back on by 3:13 PM. The MS6LCW water system needed water added to it. MIPP is once again taking beam.

Tev studies were finished at 3:39 PM.

Shot setup begins at 3:39 PM.

At 5:43 PM, experts completed repairs on BRF4. It is now up and running.

Operations established store 4584 at 6:05 PM with an initial luminosity of 152.00 E30

Operations put the Recycler into shot setup at 6:26 PM to stash antiprotons.

Pbar returns to stacking at 7:12 PM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 11:23 PM to stash antiprotons.

Sunday January 8

The midnight shift began with Operations monitoring store #4584, the Recycler Ring (RR) in shot setup, and with MiniBooNE and NuMI taking beam.

Pbar returns to stacking at 12:44 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 5:10 AM to stash antiprotons.

Pbar returns to stacking at 6:25 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 11:17 AM to stash antiprotons.

Pbar returns to stacking at 12:49 PM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 6:08 PM to stash antiprotons.

Pbar returns to stacking at 7:07 PM, beam also returns to MCenter.

Store 4584 was terminated at 10:08 PM.

At 10:28 PM, CDF and D0 both made accesses to their collision halls.

CDF and D0 accesses were completed at 10:55 PM.

Tevatron shot setup begins at 11:45 PM.

Monday January 9

The midnight shift began with the Tev in shot setup, and with MiniBooNE, and NuMI taking beam.

Operations put the Recycler into shot setup at 1:20 AM to stash antiprotons.

Operations established store 4586 at 1:46 AM with an initial luminosity of 160.82 E30

Pbar returns to stacking at 3:25 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 7:42 AM to stash antiprotons.

The Plan for Monday

Keep the store, stack, stash, run beam to NuMI, MiniBoone, and MCenter.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/9/2006   email Fermilab